Jim is six years more than twice Pat's age now. In 6 years, Jim will be exactly twice Pat's age. How old are Jim and Pat?
2x+6+6=2x but this doesn't work.

Answer by scott8148(6628)

it doesn't work because in 6 years Pat's age is x+6

2x+6+6=2(x+6) ... 2x+12=2x+12

there is no unique solution ... Pat can be any age and Jim is "six years more than twice Pat's age"